Grade 12th passAlgebra

Find out the square root : (2+i√(a^2-4)) ...........?...............

Biplob Mondol
8 Years agoGrade 12th pass

1 Answer

Param
8 Years ago
whenever such question arises we need to find the perfect square of certain form... below is the technique and solution how to solve such kind of questions.
L.H.S = 2+i\sqrt{(a+2)(a-2)}
 
=\frac{4+2\sqrt{(a+2)}\sqrt{(a-2)}i}{2} (trying to make perfect square)
 
=\frac{\sqrt{(a+2)}^{2}+[\sqrt{(a-2)}(i)]^{2}+2\sqrt{(a+2)}\sqrt{(a-2)}i}{2}
 
=\frac{[\sqrt{(a+2)}+(i)\sqrt{(a-2)}]^{2}}{2}
 
to find the square root of this==>
 
=\sqrt{}\frac{[\sqrt{(a+2)}+(i)\sqrt{(a-2)}]^{2}}{2}
 
\frac{[\sqrt{(a+2)}+(i)\sqrt{(a-2)}]}{\sqrt{2}} A.N.S